# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ Appendix A.7 Some theorems relevant to all-qubit fixing
$\boxed{l_{pi}=l_{pj},\forall p\neq i,j,\ \mathrm{and}\ l_{ii}=l_{jj}.}$ (A.14)
Theorem A.7.2 indicates that the degree of vertices $i$ and $j$ must be equal. Also, for any vertex $p$ in the graph, either both $i$ and $j$ are adjacent to $p$ or not adjacent to $p$. Therefore, both $i$ and $j$ have same sets of neighbors. Also, if there is no edge between $i$ and $j$, the distance between them is $2$. Given any graph, $G$, we can construct a new graph, $H$, by adding two new vertices and a set of edges, such that $H$ will satisfy the conditions of Theorem A.7.2.
Theorem A.7.3.
@@ 3. Case 3:
$L(H)\vec{b}=(\underbrace{0,0,\ldots,0}_{n\text{ times}},k,-k)^{\mathsf{T}}=k\vec{b}.$ (A.16)
@@
The degree matrix, $D$, is a diagonal matrix with all $N$ entries taking the value (N1)(N-1), and thus the condition $l_{ii}=l_{jj}$ is satisfied. Furthermore, since a complete graph is all-to-all connected by construction, $l_{ij}=-1,\forall i\neq j$. Thus, the second condition is satisfied too. Therefore, the Laplacian of a complete graph satisfies Theorem A.7.2, and hence Theorem A.7.1.
@@
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
Figure A.3: Figures providing the edge weight analysis of the hypercube graph family for three different edge weight functions: logarithmic, given in sub-figures (a)-(c), where the edge weight function $w_{ij}=\mathrm{log}(j+5)$, linear, given in sub-figures (d)-(f) with $w_{ij}=j+1$, and polynomial, shown in sub-figures (g)-(i) for the polynomial function defined by $w_{ij}=j^{2}+1$.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
Figure A.4: Figures presenting the analysis for modified Margulis-Gabber-Galil graph family, for which each vertex is denoted by a tuple $(p,q)$. Plots (a)-(c) contain the analysis for graphs with logarithmic edge weights defined as $w_{(p,q)(r,s)}=\mathrm{log}(nr+s+1)+1$, while sub-figures (d)-(f) presents our results for linear edge weights, $w_{(p,q)(r,s)}=nr+s+1$, and (g)-(i) have the results for edge weight function $w_{(p,q)(r,s)}=(nr+s+1)^{2}$. Here, the number of vertices in the graph is $N=n^{2}$.
@@
Figure A.5: Figures representing the difference in minimum eigenvalues between two threshold choices, $10^{-6}$ (our main results) and $10^{-10}$, for those considered graph Laplacians whose minimum eigenvalues show a downward trend with system size, $N$.
@@
Figure A.6: Figures representing the difference in minimum eigenvalues between two threshold choices, $10^{-6}$ (our main results) and $10^{-10}$, for those considered graph incidence matrices whose minimum eigenvalues show a downward trend with system size, $N^{\prime}$.
@@
(m) (n) (o)
(p) (q) (r)
(s) (t) (u)
(v) (w) (x)
(y) (z) (aa)
Figure A.7: Subfigures (a)-(i) show the advantage offered by Barabási-Albert graph family for three different seed values: 10, 23 and 50. Subfigures (j)-(r) represent advantages offered by directed Gaussian random partition graph family for seed values 10, 19 and 50. From (s) - ($a$), figures demonstrate the advantages offered by planted partition graph for seed values: 10, 12 and 50.
@@
(j) (k) (l)
(m) (n) (o)
Figure A.8: Sub-figures showing $\kappa$ and $s$ behaviour of different good graphs listed in Table 1 as well as runtime ratios of various QLSs with HHL algorithm. In the sub-figure showing $\tilde{R}(n)$ behaviour, we also provide the curve $\tilde{R}(n)=n$ for reference.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
(j) (k) (l)
(m) (n) (o)
(p) (q) (r)
(s) (t) (u)
(v) (w) (x)
(y) (z) (aa)
(ab) (ac) (ad)
(ae) (af) (ag)
(ah) (ai) (aj)
(ak) (al) (am)
(an) (ao) (ap)
(aq) (ar) (as)
(at) (au) (av)
(aw) (ax) (ay)
(az) (ba) (bb)
(bc) (bd) (be)
(bf) (bg) (bh)
(bi) (bj) (bk)
(bl) (bm) (bn)
(bo) (bp) (bq)
Figure A.9: Sub-figures showing $\kappa$ and $s$ behaviour of bad graphs listed in Table 1 as well as runtime ratios of different QLSs with HHL. In the sub-figures showing $\tilde{R}(n)$ behaviour, we also provide the curve $\tilde{R}(n)=n$ for reference.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
Figure A.10: Better graph families from incidence-matrix based systems: The first, second and third column of each panel represents $\kappa$, $s$ vs. system size $N^{\prime}=N+M$, $R(N^{\prime})$ vs $N^{\prime}$ and $\tilde{R}(n)$ vs. $n$ along with a reference linear curve, $n$, for better graph listed in Table 3.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
Figure A.11: Better graph families from incidence-matrix based systems: The first, second and third column of each panel represents $\kappa$, $s$ vs. system size $N^{\prime}=N+M$, $R(N^{\prime})$ vs $N^{\prime}$ and $\tilde{R}(n)$ vs. $n$ along with a reference linear curve, $n$, for better graphs listed in Table 3.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
(j) (k) (l)
Figure A.12: Sub-figures showing $\kappa$ and $s$ behaviour of different good graphs listed in Table 3 as well as runtime ratios of different QLSs with HHL. In the third panel, we also give plot of $\tilde{R}(n)=n$ for reference.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
(j) (k) (l)
(m) (n) (o)
(p) (q) (r)
Figure A.13: Sub-figures showing $\kappa$ and $s$ behaviour of different bad graphs listed in Table 3 as well as runtime ratios of different QLSs with HHL. In the sub-figure showing $\tilde{R}(n)$ behaviour, we also provide the curve $\tilde{R}(n)=n$ for reference.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
(j) (k) (l)
(m) (n) (o)
(p) (q)
Figure A.14: We have presented one instance from those graph families whose condition number was polylog function of $\mathcal{N}$. We observe that for all these graph families, the entries in the adjacency matrix, thus the Laplacian matrix appear diffused throughout the matrix.
(a) (b) (c)
(d) (e) (f)
(g) (h) (i)
(j) (k) (l)
Figure A.15: We have presented one instance from those graph families whose condition number was a polynomial function of $\mathcal{N}$. For all the graph families mentioned here, we witness a sharp structure in their adjacency matrix and so in its Laplacian matrix.

### v4

@@ Acknowledgements.
$\boxed{l_{pi}=l_{pj},\forall p\neq i,j,\ \mathrm{and}\ l_{ii}=l_{jj}.}$ (S.14)
Theorem S.7.2 indicates that the degree of vertices $i$ and $j$ must be equal. Also, for any vertex $p$ in the graph, either both $i$ and $j$ are adjacent to $p$ or not adjacent to $p$. Therefore, both $i$ and $j$ have same sets of neighbors. Also, if there is no edge between $i$ and $j$, the distance between them is $2$. Given any graph, $G$, we can construct a new graph, $H$, by adding two new vertices and a set of edges, such that $H$ will satisfy the conditions of Theorem S.7.2.
Theorem S.7.3.
@@ 3. Case 3:
$L(H)\vec{b}=(\underbrace{0,0,\ldots,0}_{n\text{ times}},k,-k)^{\mathsf{T}}=k\vec{b}.$ (S.16)
@@
The degree matrix, $D$, is a diagonal matrix with all $N$ entries taking the value (N1)(N-1), and thus the condition $l_{ii}=l_{jj}$ is satisfied. Furthermore, since a complete graph is all-to-all connected by construction, $l_{ij}=-1,\forall i\neq j$. Thus, the second condition is satisfied too. Therefore, the Laplacian of a complete graph satisfies Theorem S.7.2, and hence Theorem S.7.1.
@@
(a) (b)
(c) (d)
(e) (f)
Figure S.3: Figures providing the edge weight analysis of the hypercube graph family for three different edge weight functions: logarithmic, given in sub-figures (a) and (b), where the edge weight function $w_{ij}=\mathrm{log}(j+5)$, linear, given in sub-figures (c) and (d) with $w_{ij}=j+1$, and polynomial, shown in sub-figures (e) and (f) for the polynomial function defined by $w_{ij}=j^{2}+1$.
(a) (b)
(c) (d)
(e) (f)
Figure S.4: Figures presenting the analysis for modified Margulis-Gabber-Galil graph family, for which each vertex is denoted by a tuple $(p,q)$. Plots (a) and (b) contain the analysis for graphs with logarithmic edge weights defined as $w_{(p,q)(r,s)}=\mathrm{log}(nr+s+1)+1$, while sub-figures (c) and (d) presents our results for linear edge weights, $w_{(p,q)(r,s)}=nr+s+1$, and (e) and (f) show the results for edge weight function $w_{(p,q)(r,s)}=(nr+s+1)^{2}$. Here, the number of vertices in the graph is $N=n^{2}$.
@@
Figure S.5: Figures representing the difference in minimum eigenvalues between two threshold choices, $10^{-6}$ (our main results) and $10^{-10}$, for those considered graph Laplacians whose minimum eigenvalues show a downward trend with system size, $N$.
@@
Figure S.6: Figures representing the difference in minimum eigenvalues between two threshold choices, $10^{-6}$ (our main results) and $10^{-10}$, for those considered graph incidence matrices whose minimum eigenvalues show a downward trend with system size, $N^{\prime}$.
(a) (b)
(c) (d)
(e) (f)
(g) (h)
(i) (j)
(k) (l)
(m) (n)
(o) (p)
(q) (r)
Figure S.7: Subfigures (a)-(f) show the advantage offered by Barabási-Albert graph family for three different seed values: 10, 23 and 50. Subfigures (g)-(l) represent advantages offered by directed Gaussian random partition graph family for seed values 10, 19 and 50. From (m) - (r), figures demonstrate the advantages offered by planted partition graph for seed values: 10, 12 and 50.
(a) (b)
(c) (b)
(e) (f)
(g) (h)
(i) (j)
Figure S.8: Sub-figures showing $\kappa$ and $s$ behaviour of good graphs listed in Table II of the main manuscript.
(a) (b)
(c) (d)
(e) (f)
(g) (h)
(i) (j)
(k) (l)
(m) (n)
(o) (p)
(q) (r)
(s) (t)
(u) (v)
(w) (x)
(y) (z)
(aa) (ab)
(ac) (ad)
(ae) (af)
(ag) (ah)
(ai) (aj)
(ak) (al)
(am) (an)
(ao) (ap)
Figure S.9: Sub-figures showing $\kappa$ and $s$ behaviour of bad graphs listed in Table III of main manuscript. We also plot the runtime ratios for different QLSs, considered in our study, with the numerical data.
(a) (b)
(c) (d)
(e) (f)
(g) (h)
(i) (j)
(k) (l)
(m) (n)
(o) (p)
(q) (r)
(s) (t)
(u) (v)
(w) (x)
Figure S.10: Sub-figures showing $\kappa$ and $s$ behaviour of different good graphs listed in Table IV of the main manuscripts as well as as well as runtime ratios of different QLSs obtained from our numerical data. Sub-figures of first two graph families from Table IV can be found in the main manuscript.
(a) (b)
(c) (d)
(e) (f)
(g) (h)
Figure S.11: Sub-figures showing $\kappa$ and $s$ behaviour of different bad graphs listed in Table V of the main manuscript and plots of runtime ratios with different QLSs considered in our study.
@@
(m) (n)
Figure S.12: We have presented one instance from those graph families whose condition number was polylog function of $\mathcal{N}$. We observe that for all these graph families, the entries in the adjacency matrix, thus the Laplacian matrix appear diffused throughout the matrix.
@@
Figure S.13: We have presented one instance from those graph families whose condition number was a polynomial function of $\mathcal{N}$. For all the graph families mentioned here, we witness a sharp structure in their adjacency matrix and so in its Laplacian matrix.